An amount of $100,000 is invested at an annual interest rate of 6.25%. If the interest is
calculated yearly, how long would it be before the investment reached $150, 000? What
is the answer if the interest is paid monthly?.
1
Expert's answer
2021-02-24T12:29:22-0500

If the interest is calculated yearly, then:

"100,000\u00d71.0625^x = 150,000,"

"\log(1. 0625) 1.5 = 6.69"

years.

If the interest is paid monthly, then:

"100,000\u00d7(1 + 0.0625\/12)^{12x} = 150,000,"

"\log(1.0052) 1.5/12 = 6.51" years.